OpenStudy (anonymous):

Please Help! Prove that the given equation is an identity. 1 + cos 2t = cot t sin 2t

OpenStudy (anonymous):

\[1+\cos^2x - \sin^2x = 2\sin{x}\cos{x} \cdot \cot{t}\]\[1 + \cos^2x - \sin^2x = 2\cos^2{x}\]\[1 - \sin^2x = cos^2x\]\[sin^2x + cos^2x = 1\]

OpenStudy (anonymous):

The first line uses double angle formulas.